# Online Convex Optimization with Memory and Limited Predictions

### v1

@@ 4 Regret Analysis
Thus, the convergence of the first term on the right-hand-side of (12) achieves the same linear convergence rate (i.e., $(1-\mu/\beta^{\prime})^{K}$) as standard first-order method based on gradient descent for smooth and strongly convex cost functions [30]. Supposing $\varphi_{t}=0$ for all $t\geq 0$, the second term on the right-hand-side of (12) can be made arbitrarily small by setting $\delta^{\prime}$. A careful inspection of the proof of Theorem 2 reveals that the result in Theorem 2 still holds if we remove Assumption 1(c). The reason is that in the proof, we only need an upper bound on $\nabla C_{T}(\bm{x})$ for all $\bm{x}\in\mathcal{X}^{dT}$. Since $C_{T}(\cdot)$ is $\beta\sqrt{2h}$-smooth from Lemma 1, i.e., $\nabla C_{T}(\cdot)$ is continuous, and $\mathcal{X}^{dT}$ is bounded by Assumption 2, such an upper bound on $\nabla C_{T}(\cdot)$ always exists. Based on the above arguments, under the smoothness and strong convexity assumption and setting $\varphi_{t}=0$ for all $t\geq 0$, our convergence rate given in Theorem 2 for the zeroth-order subroutine in Algorithm 1 improves that of the zeroth-order method provided in [28]. [^4] Since [28] considers the setting of $\varphi_{t}=0$, we also $\varphi_{t}=0$ for comparison. Specifically, the zero-order method in [28] also has the form $\bm{x}^{j+1}=\bm{x}^{j}-\alpha\bm{g}^{j}$ as we described above, where $\bm{g}^{j}$ is obtained based on function values at certain points perturbed by a Gaussian smoothing, i.e., $u_{s}^{j+1}$ in line 14 of Algorithm 1 is drawn from a Gaussian distribution $\mathcal{N}(0,I_{d})$. Setting the step size to satisfy $\alpha=\frac{1}{4(T+4)\beta^{\prime}}$, where $Td$ is the dimension of the variable $\bm{x}^{j}$, [28] provides a convergence
$\mathbb{E}\big{[}C_{T}(\bm{x}^{K})-C_{T}(\bm{x}^{\star})\big{]}\leq\Big{(}1-\frac{\mu}{4(T+4)\beta^{\prime}}\Big{)}^{K}\big{(}C_{T}(\bm{x}^{0})-C_{T}(\bm{x}^{\star})\big{)}+\mathcal{O}(\frac{\delta^{\prime}T^{2}d^{2}\beta^{\prime}}{\mu}).$ (13)
@@
Overall Bound on the Regret of Algorithm 1. Combining the result in Theorems 1-2 gives the desired dynamic regret of Algorithm 1 for solving OCO with memory and limited predictions.
@@ 5 Numerical Experiments
This section provides numerical results to complement our theoretical analysis. Recalling that Algorithm 1 uses two-point feedback via the simulator $l_{t}(\cdot)$ and receives the predictions, for comparison, we also consider our algorithm in the single-point feedback case which has also been considered in the classic bandit OCO framework [5, 38]. Specifically, in lines 4 and 5 in Algorithm 1, we set $\bar{x}_{s}^{0}=x_{s}^{0}+\delta u^{0}$ and receive the prediction $l_{s}(\bar{x}^{0}_{s-h+1:s})$. Next, set $g_{s}^{0}=\frac{1}{\delta}\big{[}l_{s}(\bar{x}_{s-h+1:s}^{0})\big{]}u^{0}$ and then update in the same way as line 7. Similarly, for lines 9-20 of Algorithm 1, we modify the algorithm into the case of single-point feedback. Note that in our experiments we simply set $\varphi_{t}=0$ in Assumption 3, i.e., $l_{t}(\bm{x})=f_{t}(\bm{x})$ for all ${\bm{x}}\in\mathcal{X}^{d}$ and all $t\geq 0$.
Now we consider an unconstrained quadratic programming problem given by $\min_{x}C(x)=\frac{1}{2}x^{\top}Px+q^{\top}x$. Let $C(x)=\sum_{t=1}^{T}f_{t}(x_{t-h+1:t})$ and $f_{t}(x_{t-h+1:t})=\frac{1}{2}x_{t-h+1:t}^{\top}A_{t}x_{t-h+1:t}+B_{t}^{\top}x_{t-h+1:t}$. In our experiments, we consider $x_{t}\in\mathbb{R}$ and set $h=2$ and then randomly generate positive definite matrices $A_{t}$ and $B_{t}$ to make sure that Assumption 1(a) and (b) is true. Next, we set $x_{0}=0.5,\eta_{t}=0.2/t,\alpha=0.05,\delta=0.2,\delta^{\prime}=0.0001$ which satisfy the requirements stated in Theorems 1-2. Then, we consider the following three distributions for smoothing. (a) Gaussian distribution: In line 3 and line 14 of Algorithm 1, we instead pick a random variable $u\sim{\mathcal{N}}(0,1)$. (b) Truncated Gaussian distribution: To obtain the simulation results, we slightly modify Definition 1 and pick a truncated Gaussian from the interval $[-2,2]$ in lines 4 and 16 of Algorithm 1. (c) Bernoulli distribution: Note that for scalar $x_{t}\in\mathbb{R}$, the method of smoothing over a sphere is equivalent to that the random variable $u$ is drawn from a Bernoulli distribution with a $\frac{1}{2}$ probability of getting $1$ or $-1$. Hence, we pick a random variable $u$ from the Bernoulli in line 3 and line 14 of Algorithm 1. The results obtained below for Gaussian smoothing and Bernoulli smoothing are averaged over $200$ independent experiments, and the results for truncated Gaussian smoothing are averaged over $50$ independent experiments. The shaded areas in all the plots below represent quartiles.
(a) a
(b) c
(c) e
(d) b
(e) d
(f) f
Figure 1: Regret results for the initialization phase of Algorithm 1.
To verify Theorem 1, we gradually increase $T$ from $5$ to $20$ to get plots of $Reg/\sqrt{T}$, $Reg/T$ with $T$ when the random variable used for smoothing is drawn from Gaussian distribution and truncated Gaussian distribution and in both the single-point feedback and two-point feedback cases (See Figure 1). The $Reg$ here represents the regret corresponding to the initialization phase of Algorithm 1, i.e., when OCO is used in initialization. Since the regret bound in Theorem 1 scales as $\tilde{\mathcal{O}}(\sqrt{T}\ln{T})$, we first divide the regret value by $\sqrt{T}$ to get a curve that approximates the log function and then, we divide the regret value by $T$ to get a declining curve. By comparing these figures, it can be seen that the result of two-point feedback is better than that of single-point feedback, and the result when the random variable is drawn from the truncated Gaussian distribution is better than that from the Gaussian distribution. Moreover, the noise fluctuation in the Gaussian distribution is much larger than that in the truncated Gaussian distribution. Finally, the curve in Figure 1(e) changes very slowly, and roughly aligns with the curve of the logarithmic function $\ln{T}$. Figure 1(f) shows a declining curve that satisfies the characteristics of the function $\ln{T}/\sqrt{T}$. Hence, our experimental results are consistent with the theoretical analysis.
(a) a
(b) c
(c) e
(d) b
(e) d
(f) f
Figure 2: Overall regret results of Algorithm 1.
To verify Theorem 2, we set $T=20$ and $W$ is gradually increased to obtain the results that are plots of the value $\log{(Reg)}$ with $W$ when the random variable used for smoothing is drawn from the three distributions described above and in both the single-point feedback and two-point feedback settings. The $Reg$ here represents the overall regret of Algorithm 1 which consists of the two phases. As can be seen from Figure 2, the performance of two-point feedback is better than that of single-point feedback in general, and the performance of truncated Gaussian distribution is better than that of Gaussian distribution. According to Figure 2(e), the regret of our algorithm decays approximately linearly on a log scale, indicating the exponential decay with $W$. Comparing Figure 2(d) and Figure 2(e), truncated Gaussian distribution achieves better regret value and faster convergence rate than Gaussian distribution as we discussed at the end of Section 4, and all these above are consistent with the theoretical analysis of the algorithm. Finally, the performance of using the Bernoulli random variable for smoothing is comparable to the case of using the truncated Gaussian random variable, however, there is no theoretical guarantee of the regret for Bernoulli smoothing which is left to future work.
@@ A.1 Proof of Lemma 1
To relate $\nabla f_{t}(x_{t-h+1:t})$ to $\nabla C_{T}(\bm{x})$, we first augment the gradient vector $\nabla f_{t}(x_{t-h+1:t})\in\mathbb{R}^{dh}$ to a higher dimensional vector ~$\tilde{\nabla}f_{t}(x_{t-h+1:t})\in\mathbb{R}^{dT}$ by adding zero entries. By the definition of $C_{T}(\cdot)$, we see that $\nabla C_{T}(\bm{x})=\sum_{t=1}^{T}\tilde{\nabla}f_{t}(x_{t-h+1:t})$. Summing (A.5) over $t\in[T]$ now gives us
@@
which proves that $C_{T}(\cdot)$ is $\mu$-strongly convex. To show the smoothness of $C_{T}(\cdot)$, we again start from $f_{t}(\cdot)$ and its $\beta$-smoothness gives us [^5] It is well known that a $\beta$-smooth function $f:\mathbb{R}^{d}\to\mathbb{R}$ as per definition 2 satisfies that $f(y)-f(x)\leq\left\langle\nabla f(x),y-x\right\rangle+\frac{\beta}{2}\left\|y-x\right\|^{2},\ \forall x,y\in\mathbb{R}^{d}$ [31].
@@ A.2 Proof of Lemma 2
Proof of part (a) Eq. (6) follows directly from [28, Eq. (26)]. Since $f_{t}^{c}(\cdot)$ is $\beta$-smooth by Assumption 1(a), Eq. (7) follows from [28, Theorem 1].
@@
=$\overset{(a)}{=}\mathbb{E}_{\bm{v}}\Big{[}\sum_{t=1}^{T}\frac{\partial f_{t}(\bar{x}_{t-h+1:t})}{\partial x_{s}}\Big{]}$
@@
=$\overset{(b)}{=}\frac{\partial\mathbb{E}_{\bm{v}}\big{[}\sum_{k=s}^{s+h-1}f_{k}(\bar{x}_{k-h+1:k})\big{]}}{\partial x_{s}},$ (A.9)
where we can swap the order of $\mathbb{E}_{\bm{v}}$ and $\frac{\partial}{\partial x_{s}}$ in (a) and (b) since $\frac{\partial f_{t}(\bar{x}_{t-h+1:t})}{\partial x_{s}}$ is continuous in $x_{s}$ for all $t\in[T]$. Now, we notice that $\sum_{k=s}^{s+h-1}f_{k}(x_{k-h+1:k})$ may be viewed as a function of $x_{s-h+1},\dots,x_{s+h-1}$, denoted as $F_{s}(x_{s-h+1:s+h-1})$. Since $\bm{v}=(v_{1},\dots,v_{T})\sim\widetilde{\mathcal{N}}(0,I_{dT})$, we have $v_{t}\sim\widetilde{\mathcal{N}}(0,I_{d})$ for all $t\in[T]$. Letting $\bm{v}_{s}=v_{s-h+1:s+h-1}\in\mathbb{R}^{2h-1}$ be a column vector and also recalling that $\bar{x}_{t}=x_{t}+\delta^{\prime}v_{t}$ and $\tilde{x}_{t}=x_{t}-\delta^{\prime}v_{t}$, we again have from [28, Eq. (26)] that
@@ B.1 Proof of Theorem 1
Proof.
Based on the definition of Algorithm 1, let us further define an auxiliary function
@@
where $(a)$ follows from the $G$-Lipschitz of $f_{t}(\cdot)$ in Assumption 1(c) and the definition of $f_{t}^{c}(\cdot)$.
Let us denote $e_{t}=e_{t}^{m}+e_{t}^{c}$ with
@@
The update rule in line 7 of Algorithm 1 can then be viewed as performing online gradient descent on the function sequence $h_{1}(\cdot),h_{2}(\cdot),\dots$ with $x_{t+1}^{0}=\mathbf{\Pi}_{\mathcal{X}}(x_{t}^{0}-\eta_{t}(\nabla h_{t}(x_{t}^{0})+e_{t}))$, where $e_{t}$ is an additive error to the true gradient $\nabla h_{t}(x_{t})=\tilde{g}_{t}$. Based on this observation, we show in Appendix C that the following inequality holds:
$\sum_{t=1}^{T}\big{(}h_{t}(x_{t}^{0})-$ $h_{t}(x_{t}^{\star})\big{)}\leq\sum_{t=1}^{T-1}\frac{D\left\|x_{t}^{\star}-x_{t+1}^{\star}\right\|}{\eta_{t}}+\sum_{t=1}^{T}D\left\|e_{t}\right\|+\sum_{t=1}^{T}\eta_{t}\left\|\nabla h_{t}(x_{t}^{0})\right\|^{2}+\sum_{t=1}^{T}\eta_{t}\left\|e_{t}\right\|^{2}.$ (B.4)
Bounding (the expectation of) each term on the right-hand-side of (B.4) and noting that $\mathbb{E}[h_{t}(x_{t}^{0})]=\mathbb{E}[\hat{f}_{t}^{c}(x_{t}^{0})]$ and $\mathbb{E}[h_{t}(x_{t}^{\star})]=\mathbb{E}[\hat{f}_{t}^{c}(x_{t}^{\star})]$ as we argued above, we show that
$\mathbb{E}$ $\Big{[}\sum_{t=1}^{T}\hat{f}_{t}^{c}(x_{t}^{0})-\hat{f}_{t}^{c}(x_{t}^{\star})\Big{]}\leq\frac{\sqrt{2}D}{\mu}\sum_{t=1}^{T-1}\mathbb{E}\Big{[}\left\|x_{t}^{\star}-x_{t+1}^{\star}\right\|\Big{]}+(\frac{8G^{2}h^{2}+hG^{2}}{2\mu(2h-1)}+\frac{h^{3}G^{2}D}{\delta(2(2h-1))^{1/2}})(1+\ln T)$
$+\frac{2}{\delta^{2}\mu\sqrt{2h-1}}\sum_{t=1}^{T}\varphi_{t}^{2}+(\frac{D}{\delta(2(2h-1))^{1/4}}+\frac{\sqrt{2}h^{2}GD}{2\delta(2(2h-1))^{1/2}})\sum_{t=1}^{T}\varphi_{t}.$ (B.5)
The proof of (B.5) is also included in Appendix C.
Now, we can relate $\hat{f}_{t}^{c}(\cdot)$ to $f_{t}^{c}(\cdot)$ using Lemma 2. Specifically, we have that for any $t\in[T]$,
$|f_{t}^{c}(x_{t}^{0})-\hat{f}_{t}^{c}(x_{t}^{0})|\leq\frac{\delta^{2}\beta d}{2},\ |f_{t}^{c}(x_{t}^{\star})-\hat{f}_{t}^{c}(x_{t}^{\star})|\leq\frac{\delta^{2}\beta d}{2}.$ (B.6)
To complete the proof of the theorem, we relate $f_{t}^{c}(\cdot)$ to $f_{t}(\cdot)$. From the Lipschitz assumption on $f_{t}(\cdot)$, one can show that
@@
$\leq G\sum_{t=1}^{T}\sum_{j=1}^{h-1}\mathbb{E}\big{[}\left\|x_{t}^{0}-x_{t-j}^{0}\right\|\big{]}$
$\leq G\sum_{t=1}^{T}\sum_{j=1}^{h-1}\sum_{l=1}^{j}\mathbb{E}\big{[}\left\|x_{t-l+1}^{0}-x_{t-l}^{0}\right\|\big{]}$

### v3

@@ 4 Regret Analysis
Moreover, the convergence of the first term on the right-hand side of (12) achieves the same linear convergence rate (i.e., $(1-\mu/\beta^{\prime})^{K}$) as the standard first-order method based on gradient descent for smooth and strongly convex cost functions nesterov2013introductory. Supposing $\varphi_{t}=0$ for all $t\geq 0$, the second term on the right-hand side of (12) can be made arbitrarily small by setting $\delta^{\prime}$. A careful inspection of the proof of Theorem 2 reveals that the result in Theorem 2 still holds if we remove Assumption 1(c). The reason is that in the proof, we only need an upper bound on $\nabla C_{T}(\bm{x})$ for all $\bm{x}\in\mathcal{X}^{dT}$. Since $C_{T}(\cdot)$ is $\beta\sqrt{2h}$-smooth from Lemma 1, i.e., $\nabla C_{T}(\cdot)$ is continuous, and $\mathcal{X}^{dT}$ is bounded by Assumption 2, such an upper bound on $\nabla C_{T}(\cdot)$ always exists. Based on the above arguments, under the smoothness and strong convexity assumption and setting $\varphi_{t}=0$ for all $t\geq 0$, our convergence rate given in Theorem 2 for the zeroth-order subroutine in Algorithm 1 improves that of the zeroth-order method provided in nesterov2017random. [^4] Since nesterov2017random considers the setting of $\varphi_{t}=0$, we also $\varphi_{t}=0$ for comparison. Specifically, the zeroth-order method in nesterov2017random also has the form $\bm{x}^{j+1}=\bm{x}^{j}-\alpha\bm{g}^{j}$ as we described above, where $\bm{g}^{j}$ is obtained based on function values at certain points perturbed by a Gaussian smoothing, i.e., $u_{s}^{j+1}$ in line 12 of Algorithm 1 is drawn from a Gaussian distribution $\mathcal{N}(0,I_{d})$. Setting the step size to satisfy $\alpha=\frac{1}{4(Td+4)\beta^{\prime}}$, where $Td$ is the dimension of the variable $\bm{x}^{j}$, nesterov2017random provides a convergence
$\mathbb{E}\big[C_{T}(\bm{x}^{K})-C_{T}(\bm{x}^{\star})\big]\leq\Big(1-\frac{\mu}{8(Td+4)\beta^{\prime}})\Big)^{K}\big(C_{T}(\bm{x}^{0})-C_{T}(\bm{x}^{\star})\big)+\mathcal{O}(\frac{\delta^{\prime}T^{2}d^{2}\beta^{\prime}}{\mu}).$ (13)
@@
Furthermore, the zeroth-order method based on Gaussian smoothing considered in nesterov2017random provides no guarantee that its perturbed query points (e.g., $\bar{x}=x+\delta u$ with $u\sim\mathcal{N}(0,I_{d})$) remain within the feasible set $\mathcal{X}$, due to the fact that the norm of the Gaussian random vector $u$ may not be bounded. In contrast, as we argued in Remark 1, we can guarantee that the perturbed query points (e.g., $\bar{x}=x+\delta u$ with $u\sim\tilde{N}(0,I_{d})$) remain within the feasible set $\mathcal{X}$, by projecting $x$ onto smaller set $(1-\xi)\mathcal{X}$ and leverging the fact that the truncated Gaussian random vector $u$ always has bounded norm.
Overall Bound on the Regret $R(\mathcal{A})$ of Algorithm 1. Combining the result in Theorems 1-2 gives the desired dynamic regret of Algorithm 1 for solving OCO with memory and limited predictions. The total regret consists of an exponentially decaying term of the initialization regret with respect to $K=\lfloor W/(h-1)\rfloor$, plus the error terms $\varepsilon/\gamma$ and $\xi^{\prime}GT\sqrt{h}D$. Supposing $\varphi_{t}=0$ for all $t\geq 0$, the two error terms can be made arbitrarily small by setting $\delta^{\prime}$, and the overall regret Algorithm 1 reduces to the exponential decay term of the initialization regret, which matches that of the algorithm proposed in li2020online for the full information setting with one-step memory (i.e., $h=1$).
@@ 5 Numerical Experiments
This section provides numerical results to complement our theoretical analysis and demonstrate the performance of our proposed algorithm.
5.1 Experimental Setup
We consider an unconstrained quadratic programming problem, a standard benchmark in convex optimization, where the total cost is given by $C(x)=\sum_{t=1}^{T}f_{t}(x_{t-h+1:t})$. The cost at each time step is $f_{t}(x_{t-h+1:t})=\frac{1}{2}x_{t-h+1:t}^{\top}A_{t}x_{t-h+1:t}+B_{t}^{\top}x_{t-h+1:t}$. For all experiments, we consider a scalar decision variable $x_{t}\in\mathbb{R}$ (i.e., $d=1$) and a memory of $h=2$. The matrices $A_{t}\in\mathbb{R}^{2\times 2}$ and vectors $B_{t}\in\mathbb{R}^{2}$ are randomly generated at each run, with $A_{t}$ being constructed as positive definite to ensure that the strong convexity and smoothness properties from Assumption 2 hold. The algorithm parameters are set according to our theoretical analysis: $x_{0}=0.5,\eta_{t}=0.2/t,\alpha=0.05,\delta=0.2,\delta^{\prime}=0.0001$. We assume a perfect simulator, i.e., $l_{t}(x)=f_{t}(x)$ for all $t$ ($\phi_{t}=0$). We evaluate the performance of our algorithm under three different smoothing distributions used for the two-point feedback gradient estimation:
  • (a)
    Gaussian distribution: The random direction $u$ is drawn from a standard normal distribution, $u\sim\mathcal{N}(0,1)$.
  • (b)
    Truncated Gaussian distribution: The random direction $u$ is drawn from a normal distribution truncated to the interval [2,2][-2,2], which reduces the variance of the gradient estimate.
  • (c)
    Bernoulli distribution: The random direction $u$ is drawn from a Bernoulli distribution, taking values in $\{-1,1\}$ with equal probability.
All results are generated by averaging over 100 independent experimental runs. The plots show the median regret, with the shaded areas representing the interquartile range (from the 25th to the 75th percentile).
5.2 Validation of Initialization Phase Regret (Theorem 1)
First, we empirically validate the performance of the initialization phase of our algorithm (lines 3-8 in Algorithm 1), which is an Online Convex Optimization (OCO) procedure with memory. Theorem 1 establishes a dynamic regret bound of $\mathcal{O}(TlnT)$ for this phase. To verify this, we run the initialization phase over a long time horizon, varying $T$ from 5 to 500, and compute the regret of the generated sequence $x_{0}$. Figure 1 displays the normalized regret for all three smoothing distributions. The subplots show that the regret grows sub-linearly with $T$, as evidenced by the decreasing curves in the $Reg/T$ plots. Furthermore, the $Reg/T$ plots show a slowly growing trend, consistent with the theoretical $\ln T$ factor. These results provide strong empirical support for the regret bound established in Theorem 1. Among the smoothing methods, the lower-variance distributions (Truncated Gaussian and Bernoulli) tend to exhibit slightly better and more stable performance than the standard Gaussian.
(a) Gaussian
(b) Truncated Gaussian
(c) Bernoulli
Figure 1:
5.3 Regret Decay with Prediction Window (Theorem 2)
Next, we investigate the core claim of our paper: the total regret of the full algorithm decays exponentially with the length of the prediction window, $W$. This behavior is a direct consequence of the linear convergence of the zeroth-order phase, as analyzed in Theorem 2. For this experiment, we fix the time horizon at $T=400$ and vary the prediction window $W$ from 4 to 50. The regret is calculated based on the final output sequence $x_{K}$, where $K=\lfloor W/(h-1)\rfloor$ is the number of zeroth-order iterations. Figure 2 plots the natural logarithm of the regret, $\ln(Regret)$, against $W$. As depicted, all three smoothing methods show a clear, approximately linear decrease on this semi-log plot. This linear trend provides strong empirical evidence of the exponential decay of regret with respect to the prediction window size. This result validates the convergence analysis of Theorem 2, demonstrating that as more future information is made available (larger $W$), the algorithm performs more optimization iterations and rapidly converges to a better solution, drastically reducing the overall regret. Finally, the performances of using the Bernoulli random variable and the Gaussian random variable for smoothing are comparable to the case of using the truncated Gaussian random variable; however, there are no theoretical guarantees of the regret for Bernoulli smoothing or Gaussian smoothing, which are left to future work.
Figure 2: Log of regret for the full algorithm versus the prediction window size $W$, with $T=400$. The linear downward trend for all smoothing methods empirically confirms the theoretical exponential decay of regret, a key result derived from Theorem 2.
5.4 Convergence Rate Comparison with nesterov2017random
Finally, to highlight the advantages of our zeroth-order method based on truncated Gaussian smoothing, we directly compare its convergence rate against the zeroth-order method proposed by nesterov2017random based on Gaussian smoothing. In this experiment, we fix the horizon to be $T=200$. Both algorithms start from the same initial sequence $x_{0}$, generated by our OCO initialization phase. Figure 3 aggregates the performance curves under different step-size settings into a single plot for direct comparison. As we argued in Section 4, our algorithm allows for a large step size $\alpha=1/\beta^{\prime}$ compared to the step size $\alpha=1/(T\beta^{\prime})$ (here $\beta^{\prime}=20$ according to the setup in Section 5.1) that can be chosen for the algorithm in nesterov2017random, which aligns with the results in Figure 3. Moreover, when choosing the step size to be $\alpha=1/\beta^{\prime}$ for the algorithm in nesterov2017random, we see from Figure 3 that the algorithm does not converge. This divergence occurs because the method in nesterov2017random theoretically requires a step size scaling as $\mathcal{O}(1/T)$ to ensure stability, whereas our method robustly supports a much larger step size (scaling as $\mathcal{O}(1)$). These results empirically validate that our approach not only achieves faster convergence but also offers better stability under a more aggressive step size.
Figure 3: Convergence Rate Comparison (T=200)
@@ A.1 Proof of Lemma 1
To relate $\nabla f_{t}(x_{t-h+1:t})$ to $\nabla C_{T}(\bm{x})$, we first augment the gradient vector $\nabla f_{t}(x_{t-h+1:t})\in\mathbb{R}^{dh}$ to a higher dimensional vector $\tilde{\nabla}f_{t}(x_{t-h+1:t})\in\mathbb{R}^{dT}$ by adding zero entries. By the definition of $C_{T}(\cdot)$, we see that $\nabla C_{T}(\bm{x})=\sum_{t=1}^{T}\tilde{\nabla}f_{t}(x_{t-h+1:t})$. Summing (A.5) over $t\in[T]$ now gives us
@@
which proves that $C_{T}(\cdot)$ is $\mu$-strongly convex. To show the smoothness of $C_{T}(\cdot)$, we again start from $f_{t}(\cdot)$ and its $\beta$-smoothness gives us [^5] It is well known that a $\beta$-smooth function $f:\mathbb{R}^{d}\to\mathbb{R}$ as per definition 2 satisfies that $f(y)-f(x)\leq\left\langle\nabla f(x),y-x\right\rangle+\frac{\beta}{2}\left\|y-x\right\|^{2},\ \forall x,y\in\mathbb{R}^{d}$ [bubeck2015convex].
@@ A.2 Proof of Lemma 2
Proof of part (a) Eq. (6) follows directly from [nesterov2017random, Eq. (26)]. Since $f_{t}^{c}(\cdot)$ is $\beta$-smooth by Assumption 1(a), Eq. (7) follows from [nesterov2017random, Theorem 1].
@@
$\overset{(a)}{=}\mathbb{E}_{\bm{v}}\Big[\sum_{t=1}^{T}\frac{\partial f_{t}(\bar{x}_{t-h+1:t})}{\partial x_{s}}\Big]$
@@
$\overset{(b)}{=}\frac{\partial\mathbb{E}_{\bm{v}}\big[\sum_{k=s}^{s+h-1}f_{k}(\bar{x}_{k-h+1:k})\big]}{\partial x_{s}},$ (A.9)
where we can swap the order of $\mathbb{E}_{\bm{v}}$ and $\frac{\partial}{\partial x_{s}}$ in (a) and (b) since $\frac{\partial f_{t}(\bar{x}_{t-h+1:t})}{\partial x_{s}}$ is continuous in $x_{s}$ for all $t\in[T]$. Now, we notice that $\sum_{k=s}^{s+h-1}f_{k}(x_{k-h+1:k})$ may be viewed as a function of $x_{s-h+1},\dots,x_{s+h-1}$, denoted as $F_{s}(x_{s-h+1:s+h-1})$. Since $\bm{v}=(v_{1},\dots,v_{T})\sim\widetilde{\mathcal{N}}(0,I_{dT})$, we have $v_{t}\sim\widetilde{\mathcal{N}}(0,I_{d})$ for all $t\in[T]$. Letting $\bm{v}_{s}=v_{s-h+1:s+h-1}\in\mathbb{R}^{2h-1}$ be a column vector and also recalling that $\bar{x}_{t}=x_{t}+\delta^{\prime}v_{t}$ and $\tilde{x}_{t}=x_{t}-\delta^{\prime}v_{t}$, we again have from [nesterov2017random, Eq. (26)] that
@@ B.1 Proof of Theorem 1
As outlined in the main text, the regret of the initialization phase, $R(\mathcal{A}_{\text{init}})$, is defined against the true optimal solution $\bm{x}^{\star}=(x_{1}^{\star},\dots,x_{T}^{\star})$. To avoid dependencies on the specific optimum of the constrained set $(1-\xi)\mathcal{X}$, we construct a comparison sequence $\hat{\bm{x}}^{\star}=(\hat{x}_{1}^{\star},\dots,\hat{x}_{T}^{\star})$ by scaling the global optimal solution:
$\hat{x}_{t}^{\star}=(1-\xi)x_{t}^{\star},\quad\forall t\in[T].$ (B.1)
Since we assume the feasible set $\mathcal{X}$ contains the origin (or is centered such that this scaling is valid), and $\mathcal{X}$ is convex, it holds that $\hat{x}_{t}^{\star}\in(1-\xi)\mathcal{X}$. We decompose the total regret as follows:
$\mathbb{E}[R(\mathcal{A}_{\text{init}})]$ $=\mathbb{E}\Big[C_{T}(\bm{x}^{0})-C_{T}(\bm{x}^{\star})\Big]$
[…]
We will now bound each part separately.
Part 1: Bounding the Regret against $\hat{\bm{x}}^{\star}$. This part bounds the term $\mathbb{E}\big[C_{T}(\bm{x}^{0})-C_{T}(\hat{\bm{x}}^{\star})\big]$. The proof proceeds in several steps.
step 1: Preparatory Terminology. Based on the definition of Algorithm 1, let us further define an auxiliary function
@@
where $(a)$ follows from the $G$-Lipschitz of $f_{t}(\cdot)$ in Assumption 1(c) and the definition of $f_{t}^{c}(\cdot)$. Let us denote $e_{t}=e_{t}^{m}+e_{t}^{c}$ with
@@
Step 2: Relate Total Cost $C_{T}(\cdot)$ to Unitary Cost $f_{t}^{c}(\cdot)$. First, we relate the algorithmic regret in terms of $C_{T}(\cdot)$ to the regret in terms of the $h$-unitary function $f_{t}^{c}(x)=f_{t}(x,\dots,x)$.
$C_{T}(\bm{x}^{0})-C_{T}(\hat{\bm{x}}^{\star})=\sum_{t=1}^{T}\Big(f_{t}(x_{t-h+1:t}^{0})-f_{t}(\hat{x}_{t-h+1:t}^{\star})\Big).$
We can rewrite this as:
$\sum_{t=1}^{T}\Big(f_{t}^{c}(x_{t}^{0})-f_{t}^{c}(\hat{x}_{t}^{\star})\Big)+\sum_{t=1}^{T}\Big(f_{t}(x_{t-h+1:t}^{0})-f_{t}^{c}(x_{t}^{0})\Big)-\sum_{t=1}^{T}\Big(f_{t}(\hat{x}_{t-h+1:t}^{\star})-f_{t}^{c}(\hat{x}_{t}^{\star})\Big).$
We bound the expectation of the second term using the Lipschitz property (Assumption 1(c)):
@@
$\leq G\sum_{t=1}^{T}\sum_{j=1}^{h-1}\mathbb{E}\big[\left\|x_{t}^{0}-x_{t-j}^{0}\right\|\big]\leq G\sum_{t=1}^{T}\sum_{j=1}^{h-1}\sum_{l=1}^{j}\mathbb{E}\big[\left\|x_{t-l+1}^{0}-x_{t-l}^{0}\right\|\big]$